# Think Outside the Data: Colonial Biases and Systemic Issues in Automated Moderation Pipelines for Low-Resource Languages

### v2

@@ Abstract
Most social media users come from non-English speaking countries in the Global South, where much of harmful content appears in local languages. Yet, current AI-driven moderation systems struggle with low-resource languages spoken in these regions. This work examines the systemic challenges in building automated moderation tools for these languages. We conducted semi-structured interviews with 22 AI experts working on detecting harmful content in four low-resource languages: Tamil (South Asia), Swahili (East Africa), Maghrebi Arabic (North Africa), and Quechua (South America). Our findings show that beyond the well-known data scarcity in local languages, technical issues—such as outdated machine translation systems, sentiment and toxicity models grounded in Western values, and unreliable language detection technologies—undermine moderation efforts. Even with more data, current language models and preprocessing pipelines—primarily designed for English—struggle with the morphological richness, linguistic complexity, and code-mixing. As a result, automated moderation in Tamil, Swahili, Arabic, and Quechua remains fraught with inaccuracies and blind spots. Based on our findings, we argue that these limitations are not just technical gaps but reflect deeper structural inequities that continue to reproduce historical power imbalances. We conclude by discussing multi-stakeholder approaches to improve automated moderation for low-resource languages.
@@ 1 Introduction
The largest and fastest-growing user bases of social media companies come from the Global South, where billions of users generate content in their local languages. This growth has fueled a surge in non-English harmful content—such as misinformation, hate speech, and incitement to violence—contributing to severe human rights violations across the region (Samuels 2020; Milmo 2021; Yibeltal and Muia 2023). Yet, tech companies tend to prioritize moderation for English-speaking users in the West (Legon and Alsalman 2020; Popli 2021), leaving harmful content in languages spoken in the Global South largely unchecked. This neglect has deepened social harms (Nigatu and Raji 2024) and political divides across the Global South (Samuels 2020; Milmo 2021). Simultaneously, flawed moderation systems often misclassify and remove benign content in these languages, silencing marginalized voices and restricting freedom of expression (Elswah 2024b).
Historically, many languages spoken in the Global South are considered “low-resourced” due to the lack of high-quality datasets needed for training AI models (Rowe 2022; Nicholas and Bhatia 2023; Nigatu et al. 2024), which serve as the backbone of moderation infrastructure. However, data scarcity tells only part of the story. Economic and political oppression, insufficient human expertise, and limited access to digital infrastructures further exacerbate the “low-resourcedness” of these languages (Nigatu et al. 2024). Moreover, framing the problem solely as one of data scarcity overlooks broader challenges across the moderation pipeline, such as annotation, model training, and deployment. To address this critical gap, we examine the systemic barriers hindering equitable moderation for low-resource languages and explore actionable pathways to improve these systems. Specifically, we ask:
@@
To address these questions, we conducted semi-structured interviews with 22 AI researchers and practitioners, specializing in harmful content detection and developing automated tools for diverse low-resource languages that have poor moderation support. These are: Tamil from South Asia, Swahili in East Africa, Maghrebi Arabic from North Africa, and Quechua in South America.
@@
Drawing on these findings, we use coloniality as a lens to critically examine how tech companies perpetuate digital colonialism (Kwet 2019), prioritizing profit over user safety in less profitable markets in the Global South (Nicholas and Bhatia 2023). These companies not only monopolize the data extracted from next billion users in the Global South (Coleman 2018; Couldry and Mejias 2019) but also rely on biased data sources for moderation—reinforcing harmful narratives about formerly colonized populations. We highlight how current English-centric design of one-size-fits-all moderation tools reinforces colonial impulse by ignoring the linguistic diversity of Global South languages. We argue that improving moderation for these languages requires more than technical fixes, as competing stakeholder priorities demand deeper systemic changes. The key contributions of our work are as follows:
  • A qualitative study that uses coloniality as a lens to provide a critical and nuanced understanding of how historical power imbalances disproportionately affect automated moderation pipelines for diverse low-resource languages in the Global South.
  • An outline of paths forward, acknowledging the complexity, practical constraints, and systemic issues to improve moderation for low-resource languages.
@@ 2 Related Work
In this paper, we situate our work first by discussing existing content moderation literature focusing on the Global South. We then describe scholarly work investigating colonial biases in content moderation systems.
@@ 2.1 Content Moderation in the Global South
Content moderation refers to reviewing user-generated content to see if it aligns with tech company’s policies on what content should or should not be allowed on their platforms. Most tech companies moderate content using a combination of manual human reviews and automated AI models (Gorwa 2019). However, these companies often lack financial incentives to invest in moderation resources for less profitable markets in the Global South (De Gregorio and Stremlau 2023; Nicholas and Bhatia 2023). For instance, Meta funnels 87% of its global misinformation budget to the United States (US), despite Americans comprising only 10% of its user base (Popli 2021). The disparity is even more glaring when tech companies swiftly respond to harmful content from European countries that either offer strong economic incentives (De Gregorio and Stremlau 2023) or are of geopolitical interest to the US (e.g., Russia-Ukraine war) (Meta 2022). In contrast, tech companies have been less proactive in countering disinformation campaigns and extreme speech festering in many Global South countries (Milmo 2021; Wong and Ernst 2021; Wong and Harding 2021; Yibeltal and Muia 2023), while unjustly removing culturally appropriate and politically legitimate content from this region (Elswah 2024b; Shahid and Vashistha 2023).
The inability of tech companies to accurately and fairly moderate content in the Global South is often attributed to their reliance on automated moderation systems, trained on data-rich languages like English and a handful of European languages (Nicholas and Bhatia 2023; De Gregorio and Stremlau 2023). Prior research highlights that the lack of data in low-resource languages hinders the development of robust NLP technologies for detecting harmful content in these languages (Nicholas and Bhatia 2023; Nigatu et al. 2024). In contrast, little attention is given to other critical stages of automated moderation pipelines, such as who annotates what is harmful or what assumptions are made about deploying these models in complex, low-resource environments. To address this critical gap, we examine the systemic challenges AI researchers and practitioners encounter at various stages of automated moderation pipelines when developing moderation technologies for low-resource languages in the Global South. We now present scholarly work critically examining systemic issues in content moderation systems through the lenses of power and control.
@@ 2.2 Coloniality in Content Moderation
Coloniality perpetuates historical power imbalances through extraction, enslavement, and appropriation (Mbembe 2016). Decolonial scholars argue that the colonial structures persist today by exploiting the resources and labor of historically colonized populations while reinforcing Western dominance in governance and knowledge production (Quijano 2000, 2007a). Thus, decolonial computing critically examines who participates in computing, where it occurs, and how it shapes both knowledge (epistemology) and existence (ontology) from the perspective of those at the margins of the modern world system (Ali 2016). Whereas, postcolonial scholars conceptualize coloniality in computing as when technologies designed in the West, with Western values encounter diverse cultures (Irani et al. 2010).
@@
Several scholars have critically examined content moderation systems through the lens of coloniality. For instance, Shahid and Vashistha (2023) use decoloniality as a lens to highlight how tech companies frequently impose Western values as global community standards, disregarding local socio-cultural norms when assessing online harms in the Global South. They draw parallels between Western-centrism in community guidelines and the way colonial powers systematically suppressed Indigenous and marginalized communities’ diverse ways of being, while imposing Euro-modern rationality as the only legitimate way (Said 2000; Gramsci 2020; Quijano 2007a). Similarly, Siapera (2022) provides a decolonial critique of how tech companies dismiss input from racialized users when shaping policies on racist hate speech. She argues that the company’s race-blind moderation policies mirror colonial legacies, where the identities and lived experiences of racialized communities are considered inferior (Quijano 2007b) and criminalized through “neutral” technologies (Benjamin 2023).
@@
In addition, scholars draw attention to how tech companies build AI models for moderation on the backs of low-wage moderators and marginalized communities, whose labor and trauma fuel training datasets (Siapera 2022; Shahid and Vashistha 2023; Elswah 2024a). They point out that Western tech companies treat moderators from the Global South as dispensable, often concealing the true nature of the work during recruitment, and avoiding liability for the harms these moderators experience (Ahmad and Krzywdzinski 2022; Elswah 2024a).
@@
Moreover, errors in moderation systems disproportionately affect marginalized communities, whose voices have been historically silenced. For example, AI models have been shown to drive systemic racism and heteronormative patriarchy by erroneously labeling Black and queer vernacular as toxic (Bhattacharyya 2018; Sap et al. 2019; Mohamed, Png, and Isaac 2020). Current moderation systems, shaped by a Western perspective, frequently misclassify innocuous and culturally appropriate content in non-Western contexts as harmful, while failing to detect actual harmful content (Shahid and Vashistha 2023). For example, Google’s Perspective API underestimates the toxicity of extreme speech in Swahili and Hindi, but rates similar content in Western languages, such as English and German more accurately (Udupa, Maronikolakis, and Wisiorek 2023). Udupa, Maronikolakis, and Wisiorek (2023) stressed that these errors persist because current moderation systems inherit Eurocentric colonial frameworks that rationalize uneven allocation of corporate resources for content moderation across different geographies and language communities.
@@
We contribute to this growing body of work by critically examining the often-overlooked socio-political issues embedded at different stages of automated moderation pipelines, spanning data collection, labeling, cleaning, model training, and evaluation. As most studies overlook socio-political contexts when discussing power asymmetry within the AI pipeline (Ovalle et al. 2023)—to address this gap, we interrogate the normative assumptions and design paradigms underpinning AI-driven moderation technologies through the lenses of power and control. Drawing on interviews with AI experts, we explore how prevailing language technologies shape automated moderation of online harms in diverse Global South contexts (RQ1) and how moderation practices might be improved for these languages (RQ2).
@@ 3 Methodology
To examine disparities in automated moderation pipelines, we interviewed 22 AI researchers and practitioners specializing in automatic detection of harmful content in diverse low-resource languages spoken across the Global South.

### v3

@@ Abstract
Most social media users come from the Global South, where harmful content usually appears in local languages. Yet, AI-driven moderation systems struggle with low-resource languages spoken in these regions. Through semi-structured interviews with 22 AI experts working on harmful content detection in four low-resource languages: Tamil (South Asia), Swahili (East Africa), Maghrebi Arabic (North Africa), and Quechua (South America)–we examine systemic issues in building automated moderation tools for these languages. Our findings reveal that beyond data scarcity, socio-political factors such as tech companies’ monopoly on user data and lack of investment in moderation for low-profit Global South markets exacerbate historic inequities. Even if more data were available, the English-centric and data-intensive design of language models and preprocessing techniques overlooks the need to design for morphologically complex, linguistically diverse, and code-mixed languages. We argue these limitations are not just technical gaps caused by “data scarcity” but reflect structural inequities, rooted in colonial suppression of non-Western languages. We discuss multi-stakeholder approaches to strengthen local research capacity, democratize data access, and support language-aware solutions to improve automated moderation for low-resource languages.
@@ 1 Introduction
The largest and fastest-growing social media user base comes from the Global South, where billions generate content in their local languages. This growth has fueled misinformation, hate speech, and incitement to violence in non-English languages, contributing to serious human rights violations (Samuels 2020; Milmo 2021; Yibeltal and Muia 2023). However, the moderation systems deployed by tech companies prioritize English-speaking users in the West (Legon and Alsalman 2020; Popli 2021), leaving harmful content in Global South languages largely unchecked, increasing social harm and political divides (Nigatu and Raji 2024; Samuels 2020; Milmo 2021). Moreover, flawed moderation systems often censor benign content in non-English languages and silence marginalized voices (Elswah 2024b).
Regardless of the content modality, tech companies heavily rely on text-based methods, such as analyzing captions, transcripts, and subtitles, to automatically moderate harmful content (stream 2025a, b). However, this approach poses challenges for many languages spoken in the Global South, which are historically considered “low-resourced” due to the lack of high-quality datasets needed to train AI models (Rowe 2022; Nicholas and Bhatia 2023; Nigatu et al. 2024). However, data scarcity tells only part of the story. Economic and political oppression, insufficient human expertise, and limited access to digital infrastructures further exacerbate the “low-resourcedness” of these languages (Nigatu et al. 2024). Moreover, framing the problem solely as one of data scarcity overlooks broader challenges across the moderation pipeline, such as annotation, model training, and deployment. To address this critical gap, we examine the systemic barriers hindering equitable moderation for low-resource languages and explore actionable pathways to improve these systems. Specifically, we ask:
@@
To address these questions, we conducted semi-structured interviews with 22 AI researchers and practitioners, specializing in harmful content detection and developing automated tools for diverse low-resource languages, such as Tamil from South Asia, Swahili in East Africa, Maghrebi Arabic from North Africa, and Quechua in South America.
@@
Drawing on these findings, we use coloniality as a lens to critically examine how tech companies perpetuate digital colonialism (Kwet 2019), prioritizing profit over user safety in less profitable Global South markets (Nicholas and Bhatia 2023). These companies not only monopolize data extracted from the next billion users (Coleman 2018; Couldry and Mejias 2019), but also rely on biased sources for moderation—reinforcing harmful narratives about formerly colonized populations. We highlight how the English-centric, one-size-fits-all design of moderation tools reflects a colonial impulse by ignoring the linguistic diversity of Global South languages. Improving moderation for these languages requires more than technical fixes, as competing stakeholder priorities demand deeper systemic change. The key contributions of our work are as follows:
  • A qualitative study using coloniality as a lens to examine how historical power imbalances skew automated moderation for low-resource languages in the Global South.
  • A guideline to improve moderation for low-resource languages, acknowledging the practical and systemic issues.
@@ 2.1 Content Moderation in the Global South
Content moderation refers to reviewing user-generated content to see if it aligns with tech company’s policies on what should be allowed on the platform. Most companies use a mix of manual human reviews and automated AI models (Gorwa 2019). However, they often lack financial incentives to invest in moderation resources for less profitable markets in the Global South (De Gregorio and Stremlau 2023; Nicholas and Bhatia 2023). For instance, Meta funnels 87% of its global misinformation budget to the United States (US), despite Americans comprising only 10% of its user base (Popli 2021). The disparity is more glaring when companies swiftly handle harmful content from European countries that either offer strong economic incentives or hold geopolitical interest for the US (e.g., Russia-Ukraine war) (De Gregorio and Stremlau 2023). In contrast, tech companies have been less proactive in countering harmful content festering in many Global South countries (Milmo 2021; Wong and Ernst 2021; Wong and Harding 2021; Yibeltal and Muia 2023), while unjustly removing culturally and politically legitimate content from this region (Elswah 2024b; Shahid and Vashistha 2023).
Prior research links poor moderation in the Global South to tech companies’ reliance on automated systems primarily trained on English and the lack of data in low-resource languages (Nicholas and Bhatia 2023; De Gregorio and Stremlau 2023; Nigatu et al. 2024). In contrast, little focus is given to other critical stages of moderation pipelines, such as who annotates what is harmful or what assumptions are made about deploying these models in complex, low-resource environments. To fill this critical gap, we explore the systemic challenges AI experts face throughout the moderation pipeline, placing these issues in the context of historical and colonial marginalization of non-English languages. We now present scholarly work critically examining power and control in content moderation systems.
@@ 2.2 Coloniality in Content Moderation
Decolonial scholars argue that colonial power structures persist by exploiting the labor and resources of formerly colonized populations while maintaining Western dominance over governance and knowledge production (Quijano 2000, 2007a). Decolonial computing challenges who designs technology, where it’s developed, and how it impacts marginalized communities (Ali 2016). Whereas, postcolonial scholars critique the imposition of Western-centric technologies on diverse cultural contexts (Irani et al. 2010). Scholars have applied these lenses to examine content moderation systems.
@@
Shahid and Vashistha (2023) used decoloniality to show how tech companies often impose Western values as global community standards, ignoring local socio-cultural norms when assessing online harms in the Global South. They draw parallels between Western-centrism in community guidelines and colonial suppression of Indigenous and marginalized communities’ diverse ways of being, imposing Euro-modern rationality as the only legitimate way (Said 2000; Gramsci 2020; Quijano 2007a). Similarly, Siapera (2022) offers a decolonial critique of how tech companies dismiss input from racialized users when shaping policies against racist hate speech. She argues the race-blind, “neutral” moderation technologies mirror colonial legacies, criminalizing the identities and lived experiences of racialized communities as inferior (Quijano 2007b; Benjamin 2023).
@@
In addition, scholars highlight how tech companies build AI moderation systems exploiting the labor and trauma of low-wage moderators from the Global South (Siapera 2022; Shahid and Vashistha 2023; Elswah 2024a). Tech companies often treat these moderators as disposable, concealing the nature of the work during recruitment, and evading responsibility for the harms moderators face (Ahmad and Krzywdzinski 2022; Elswah 2024a).
@@
Moreover, moderation errors disproportionately harm marginalized communities, whose voices have long been silenced. AI models often reinforce systemic racism and heteronormative patriarchy by mislabeling Black and queer vernacular as toxic (Bhattacharyya 2018; Sap et al. 2019; Mohamed, Png, and Isaac 2020). Shaped by Western norms, these systems frequently misclassify culturally appropriate content in non-Western contexts as harmful while overlooking actual harmful content (Shahid and Vashistha 2023). For instance, Google’s Perspective API underestimates toxicity in Swahili and Hindi but rates similar content in English and German more accurately (Udupa, Maronikolakis, and Wisiorek 2023). Udupa, Maronikolakis, and Wisiorek (2023) argues these biases persist because moderation systems inherit Eurocentric, colonial frameworks that justify unequal resource allocation across languages and regions.
@@
We contribute to existing literature by critically examining the often-overlooked socio-political dynamics embedded across the automated moderation pipeline—from data collection to labeling, training, and evaluation. Since many studies overlook historical contexts when discussing power asymmetries in AI (Ovalle et al. 2023), we address this gap by interrogating the normative assumptions and design paradigms behind AI-driven moderation through the lenses of power and control. Drawing on interviews with AI experts, we explore how current language technologies shape moderation in diverse Global South contexts (RQ1) and how these practices might be improved (RQ2).
@@ 3 Methodology
To examine disparities in automated moderation pipelines, we interviewed 22 AI experts specializing in diverse low-resource languages spoken across the Global South.